Tony had $9000 to invest. He invested some an account paying 8.5% simple interest and the remainder in an account paying 11% simple interest. If his total interest at the end of one year was $840, find the amount he invested in each account.

Answer by checkley79(3341)

.11X+.085(9000-X)=840
.11X+765-.085X=840
.11X-.085X=840-765
.025X=75
X=75/.025
X=$3,000 INVESTED @ 11%
9,000-3,000=$6,000 INVESTED @ 8,5%
PROOF:
.11*3,000+.085*6,000=840
330+510=840
840=840